# Bi-Directional Deep Contextual Video Compression

### v1

@@ VI Conclusion
In this paper, we propose a bi-directional deep contextual video compression scheme tailored for B-frames, termed DCVC-B. We focus on improving the compression performance for the B-frame coding from three aspects. Firstly, we propose a bi-directional motion difference context propagation method to reduce the motion coding costs. Secondly, we propose a bi-directional contextual compression model and a bi-directional temporal entropy model to make better use of multi-scale temporal contexts. Thirdly, we propose a hierarchical quality structure-based training strategy to achieve a better bit allocation within a large GOP. Experimental results show that in terms of PSNR, our DCVC-B scheme outperforms the reference software of H.265/HEVC under the random access configuration by a large margin. On some testing datasets, DCVC-B even outperforms the reference software of H.266/VVC.

### v2

@@ VII Conclusion
In this paper, we propose a bi-directional deep contextual video compression scheme tailored for B-frames, termed DCVC-B. We focus on improving the compression performance for the B-frame coding from three aspects. Firstly, we propose a bi-directional motion difference context propagation method to reduce the motion coding costs. Secondly, we propose a bi-directional contextual compression model and a bi-directional temporal entropy model to make better use of multi-scale temporal contexts. Thirdly, we propose a hierarchical quality structure-based training strategy to achieve a better bit allocation within a large GOP. Experimental results demonstrate that, in terms of PSNR, our DCVC-B scheme significantly outperforms the reference software of H.265/HEVC under the random access configuration. On some testing datasets, DCVC-B even surpasses the reference software of H.266/VVC. Additionally, we analyze the unique challenges of deep B-frame coding and identify the limitations of our scheme. To address these limitations, we propose important future work that could further enhance the compression performance of deep B-frame coding, thereby beneficial for the community.